Suppose that at a price of $500 the weekly round-trip ticket sales for a flight from New York to San Francisco are 10,000 tickets, and at a price of $1,000 the ticket sales are 9,000 tickets. Calculate the price elasticity of demand over this price range and determine if demand is elastic, inelastic, or unit elastic.
Answer #1

price elasticity of demand = (dQ / dP) * (P / Q)

dQ = 9000 - 10000 = -1000

dP = 1000 - 500 = 500

So, (-1000 / 500 ) * (500 / 10000) = -2 (500 / 10000) = - 2 * (1 / 20) = - (1/10)

So demand is inelastic.